# Register Always Matters: Analysis of LLM Pretraining Data Through the Lens of Language Variation

### v1

@@ Appendix C Numerical results
Figure 4: Individual register and HPLT v2 (deduplicated) model accuracies as a function of number of training tokens. Rolling average over 6 billion tokens applied for ease of reading overlapping lines.
Figure 5: Combination model and baseline accuracies as a function of number of training tokens. Rolling average over 6 billion tokens applied for ease of reading overlapping lines.
Figure 6: Accuracies of all models on individual benchmarks. Random-guess thresholds shown in gray when at least one model falls under it. Rolling average over 6 billion tokens applied for ease of reading overlapping lines.
@@
Table 2: Accuracies for all models, first and last checkpoint. “Accuracy” denotes accuracy weighted by benchmark, which we present in our figures. Reweighing was done to mitigate the dominance of MMLU in the results, see Section 3.3 and Appendix C. acc_norm and stderr stand for average accuracy and standard error as given by LightEval.

### v2

@@ Appendix D Numerical results
Table 3: Accuracies for all models, first and last checkpoint. “Accuracy” denotes accuracy weighted by benchmark, which we present in our figures. Reweighing was done to mitigate the dominance of MMLU in the results, see Section $3.$ acc_norm and stderr stand for average accuracy and standard error as given by LightEval.
Appendix E Performance by benchmark and register
Figures 7, 8, and 9 present the results over all checkpoints, with Figure 7 showing results by register, others by benchmark. To augment readability, we again applied a rolling average. In Figure 7, the values are normalised with respect to the average benchmark score on the final checkpoint; for example, the Description (dtp) model outperforms the average of all registers by 0.1 points on the final checkpoint on the ARC Easy benchmark.
Figure 7: Models’ performance by register. Results are scaled by the average last checkpoint performance to highlight the models’ differences. Dotted lines to increase readability.
Figure 8: Accuracies of register models on individual benchmarks. Random-guess thresholds shown in grey when at least one model falls under it. Rolling average over 6 billion tokens applied for ease of reading overlapping lines.
Figure 9: Accuracies of combination models on individual benchmarks. Random-guess thresholds shown in grey when at least one model falls under it. Rolling average over 6 billion tokens applied for ease of reading overlapping lines.